OpenStudy (anonymous):

what is the nearest point of X^4+y^4=1 to x+8y=24

OpenStudy (bahrom7893):

is this for calculus?

OpenStudy (amistre64):

you have to find some slope that are perpendicular and measure the distance between them im assuming

OpenStudy (amistre64):

lol ... well, the line only has 1 slope for every point :)

OpenStudy (bahrom7893):

use distance formula and minimize it by takin derivative..

OpenStudy (amistre64):

slope of the line = -1/8 determine when the detrivative of the other = 1/8

OpenStudy (amistre64):

...dropped me - sign :)

OpenStudy (amistre64):

Dx{x^4+y^4=1} 4x^3 + 4y^3 y' = 0 y' = (-4x^3)/(4y^3) = -x^3/y^3

OpenStudy (amistre64):

there was some sort of bacterial infestation in the ham last night and when we woke up this morning this is what we found lol

OpenStudy (amistre64):

-1 -x^3 --- = ------ ; got to test all the possibles 8 y^3 -x^3 = 1 x^3 = -1 x = -1 but then again, the negative can go south to the 8 sooo x^3 = 1 when x=1 ... what am i missing in this if anything?

OpenStudy (amistre64):

...at any rate I see combos of x=1 and y=2 ... just gotta sort out the signs :)

OpenStudy (amistre64):

we want to match these points to a line that is perp to our slope to get the shortest distances ... so the line of y = 8x -8(Px) +Py

OpenStudy (anonymous):

Maybe x is 1/2....

OpenStudy (amistre64):

\[-\left(\frac{x}{y}\right)^3=-\frac{1}{8}\] (-1,-2) and (1,2) is what I see ....

OpenStudy (anonymous):

but if those are points on the graph where the sloep of the tangent line = -1/8, i have to mention that those points aren't *on* the graph of x^4 + y^4 = 1... =/

OpenStudy (amistre64):

y1 = 8x -8(-1) -2 = 8x +6 y2 = 8x -8(1) +2 = 8x -6 perhaps?

OpenStudy (amistre64):

hmmm ... that could be an issue; but never sweat the small stuff :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

1/ 1/2 is 1/8

OpenStudy (anonymous):

Nah, can't be right..

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x^4%2By^4%3D1+and+x%2B8y%3D24 the graphs look like this

OpenStudy (amistre64):

and its the top of the skinny eclipse thats gonna be our breadwinner

OpenStudy (anonymous):

can we zoom in on it more?

OpenStudy (amistre64):

nah; but its centered at the origin

OpenStudy (amistre64):

lol ... matbe we can :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x^4%2By^4%3D1+and+x%2B8y%3D8.35 getting closer ;)

OpenStudy (anonymous):

ok...let's step back for a second... -x^3/y^3=-1/8 x^3/y^3=1/8 x/y=1/2 y=2x

OpenStudy (anonymous):

this means that the slope of the tangent line will be -1/8 when y is 2x. plug this back into the original...

OpenStudy (anonymous):

you get x= as the 4th root of 1/17

OpenStudy (anonymous):

y would then be 2 times the fourth root of 17.

OpenStudy (anonymous):

and these values work for the original equation.

OpenStudy (anonymous):

That's lookin good..

OpenStudy (anonymous):

very ugly, but i bet if you throw a line through this point with slope fo 8, it'll be a nice perpendcular to the line

OpenStudy (anonymous):

lol...i'd use point slope form on the graphing calculator =)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x^4%2By^4%3D1+and+x%2B8y%3D8.35 i got an accuracy of a few decies :)

OpenStudy (anonymous):

so now i guess we'd need to find the intersection of\[x+8y=24\]\[y-2(1/17)^{1/4}=8[x-(1/17)^{1/4}]\]

OpenStudy (anonymous):

then calculate the distance between the points... who came up with this problem?!?

OpenStudy (amistre64):

dante im guessing

OpenStudy (anonymous):

well, it doesn't seem to get any better, because the x-coordinate alone is\[x=(24-48\sqrt[4]{17})/65\] I'm not sure my brain can take anymore...

OpenStudy (anonymous):

my mistake...24+48...

OpenStudy (anonymous):

and inside the radical is 1/17..

OpenStudy (anonymous):

My 3d soft has to make this sort of calculation all of the time (collision avoidance). Have to see if I can find how they do it....